# Biased AI can Influence Political Decision-Making

### v1

@@ Abstract
As modern AI models become integral to everyday tasks, concerns about their inherent biases and their potential impact on human decision-making have emerged. While bias in models are well-documented, less is known about how these biases influence human decisions. This paper presents two interactive experiments investigating the effects of partisan bias in AI language models on political decision-making. Participants interacted freely with either a biased liberal, conservative, or unbiased control model while completing political decision-making tasks. We found that participants exposed to politically biased models were significantly more likely to adopt opinions and make decisions aligning with the AI’s bias, regardless of their personal political partisanship. However, we also discovered that prior knowledge about AI could lessen the impact of the bias, highlighting the possible importance of AI education for robust bias mitigation. Our findings not only highlight the critical effects of interacting with biased AI and its ability to impact public discourse and political conduct, but also highlights potential techniques for mitigating these risks in the future.
@@
To bridge this gap, we conducted a series of experiments to evaluate the impact of biased AI language models on human decision-making in a more typical setting, using dynamic chatbx interactions, with tasks centered on personal opinions and decisions. Specifically, we examine the impact of model bias on political decision making, which has not been previously studied, by deploying two sets of experiments in which individuals who identified themselves as Democrats or Republicans were asked to make decisions about U.S. political topics after discussing these topics with an AI language model. For this paper, we focus on language model behavioral bias, which we define as the variations in generated text, where the model’s responses—such as recognizing, rejecting, or reinforcing stereotypes—change based solely on the social group mentioned in the prompt [Kumar2024DecodingBA]. The type of model bias we examine is partisan bias, which we define as the the tendency of political partisans to process information and make judgments in a way that favors their own party [iyengar2019origins, NBERw19080].
@@
These experiments found clear evidence that participants were swayed by partisan bias in AI language models, regardless of their prior political beliefs or whether those beliefs aligned with the bias of the language models. Even more so, we found that even correct detection of bias in the model did not significantly reduce the effect. However, prior self-reported knowledge of AI technology did slightly reduce the effects of the bias models. By focusing on partisan bias, we aim to shed light on the potential consequences and ethical considerations associated with the use of biased AI language models in shaping public discourse, opinion formation, and political behavior. Further, this study is among the first to determine if dynamic interactions with AI language models lead to bias influencing human decision-making as well as investigating if these AI language models are able to lead to shifts in decision-making for decision often driven by personal values and identity.
@@ 1 Results
Participant Partisanship | Coefficient | Beta Value | t Value | p-value |
@@
Participant Partisanship | Coefficient | Beta Value | t Value | p-value |
@@
In the Topic Opinion Task, we found that participants who interacted with biased language models were more likely to change opinions in the direction of the bias of the language model compared to those who interacted with the neutral AI language model, even if it was opposite to what their beliefs were likely to be, based on their stated political affiliation. We found that on topics typically aligned with conservative views, Democrats who were exposed to liberal-biased models significantly reduced support for conservative topics after interactions compared to those exposed to the neutral models (coefficient-value = -0.85, t = -2.38, p-value = 0.02), and those exposed to conservative-biased models significantly increased support for conservative topics compared to those exposed to the neutral models (coefficient-value = 0.98, t = 2.71, p-value = .007). Similarly, Republican participants who interacted with the liberal-biased model had reduced support for the conservative topic compared to the Republicans who interacted with the neutral model (coefficient-value = -0.79, t = -2.16, p-value = .03). However, Republican participants exposed to the conservative-bias model did not have a statistically significant difference in opinions compared to those exposed to the neutral model, likely representing a ceiling effect, as these participants already agreed strongly with the model’s bias and therefore had similar opinions to when compared to the neutral AI language model. See Table 1 (top) for full results.
@@
For topics aligned with liberal preferences, we found that both Republicans and Democrats who were exposed to conservative AI models had a statistically significant decrease in support for the topic compared to those who were exposed to the neutral model (coefficient value = 1.42, t = 3.91, p-value < 0.001 and coefficient-value = 1.44, t = 3.82, p-value < 0.001, respectively). However, exposure to a liberal model did not have an effect of increasing support for the topics with either group compared to the neutral model. See Table 1 (bottom) for full results. We also conducted the same analysis subsetting only to participants who indicated no prior knowledge of the topics and the results remain unchanged, indicating that interacting with bias AI language models affects opinion formation as well (see Section E.2 for details).
@@
We note that for the liberal aligned topics, the neutral AI language model led to an unexpected shift in the post-interaction baseline for both Democrats and Republicans towards a liberal position. Thus, there was a ceiling effect, whereby exposure to a liberal-bias AI language model could not further shift an already moderate liberal opinion when interacting with the neutral model. One possible explanation for the liberal shift from the control model is that partisan respondents do not exhibit expected consistency in ideological beliefs on low salience issues with multiple dimensions [lenz2012follow, doi:10.1086/700005]. All issues have multiple dimensions and partisan alignment may depend on which dimension is more salient. Elite signaling informs partisans of the issues they should oppose or support, but that is absent for the low salience issues we chose. For more discussion, see Section E.1.
@@
Qualitatively, participants in this tasks often treated the model like a traditional search engine, with of initial interactions involving queries such as “What is <topic>?”. Common follow-up questions included “What are the pros/cons of <topic>?” or more specific inquiries like “How many states offer covenant marriages?” or “Does the US practice unilateralism in foreign relations?”. Although participants mainly sought information from the model in the form of questions, we did find that asked the model for its opinion on the topic. Another used some form of conversational language such as “hello”, “good afternoon”, “I see”, or “thank you”, suggesting that they found it to be more human-like than a simple search engine. Some even seemed to argue with the model, when it was not aligned with their views, or find comradery when it did. For examples of these conversations, see Section E.5.
@@
Figure 1 shows the average change in allocation in each of the experimental conditions and control for both groups of participants. We found that the largest average change ( confidence interval) was demonstrated for Democrat participants when exposed to the conservative AI model with average changes of () for Education, () for Welfare, () for Safety and () for Veterans. Similarly, the largest change in allocation for Republican participants was when they are exposed to the liberal AI model with average changes ( confident interval) of () for Education, () for Welfare, () for Safety, and () for Veterans. This task showed that interacting and collaborating with bias AI had strong effects on the change in outcome and final allocation of the budgets proposed.
@@
We hypothesized that prior general knowledge of AI might reduce the effects of interaction with the biased AI language model, as individuals with some understanding of AI might be more aware of the potential biases and reliability limitations of large language models. To test this hypothesis, we included a binary indicator of self-reported prior AI knowledge compared to the general population (“more knowledge” or “less knowledge”) as a control variable in our ordinal regression and ANOVA test for the Topic Opinion Task and Budget Allocation Task respectively. We note that only 32% (n=) of Democrats and 47% (n=71) of Republicans indicated having more AI knowledge. Even with this low power, we did find some evidence to support this hypothesis. Specifically, for the Topic Opinion Task, we found that for the conservatively supported topic, the effect of interacting with the AI bias was significantly reduced the effect of the bias interactions for Democrats participants who indicated having more prior knowledge compared to those who had less prior knowledge of AI (coefficient value = -0.79, t value = -2.51, p value =.01). For the Budget Allocation Task, we found significant differences at the level in the allocation between participants who reported more or less AI knowledge in the Veterans funding allocation for Democrat participants (p-value = 0.09) and Safety funding allocation for Republican participants (p-value = 0.08). See Section E.3 for the full results. These findings give some indication that prior knowledge of AI might reduce the effects of bias AI language models.
A second hypothesis, which has been supported in the context of traditional media, suggests that participants who recognize a news source’s bias are less influenced by it [kroon_etal]. Thus, we aimed to test whether awareness of the bias model itself would also alter participants’ behavior. To do this, we added a binary indicator bias detection, where “correct” detection meant that a participant was in a bias condition and correctly identified the model as biased by answering “likely yes” or “definitely yes” when asked if the model was biased. The answer was labeled as “incorrect” if it was “likely no” or “definitely no”. We note that (n=) of Democrat and 67% (n=) of Republicans correctly identifying bias in a model. Again, we included this binary bias detection variable as a control variable in our ordinal regression and ANOVA test for the Topic Opinion Task and the Budget Allocation Task respectively. For the Topic Opinion Task and Budget Allocation Task we did not find the coefficient significance in any condition. See Section E.3 for the full results. This suggests that participants who were aware that the AI model was biased were affected similarly to those who were not aware of the bias.
@@
The collaborative nature of the Budget Allocation Task provided a unique opportunity to explore the persuasion techniques used across experimental conditions, offering valuable insights for model bias mitigation. To analyze the conversations, we annotated them using the latest GPT-4 model [gpt4], employing a list of persuasion techniques compiled from a meta-analysis of persuasive strategies [Piskorski2023SemEval2023T3]. To ensure quality, we conducted a human evaluation of of the model’s annotations, achieving accuracy. Our analysis (see Figure 2(a)) found no significant differences in the distribution of persuasion techniques between the experimental conditions and the control group, as determined by a Chi-square test with Monte Carlo correction ( = 24.5, p = .07). Across all three conditions, the most frequently used techniques used by the AI language models were “Appeal to Values,” “Consequential Oversimplification,” “Appeal to Authority,” and “Repetition.”
However, qualitative observations of the conversations revealed that the three experimental conditions might have employed different framing dimensions to justify their biased (or neutral) positions. To analyze this quantitatively, we performed a similar analysis as before, using the latest GPT-4 model to annotate the Budget Allocation Task conversations with a list of framing techniques [Card2015TheMF]. Again, to validate we conducted human evaluation of of the model’s annotations, achieving accuracy. Our findings showed that the three experimental conditions employed significantly different framing dimensions, as determined by a Chi-square test with Monte Carlo correction (, p-value ). Furthermore, both the liberal and conservative bias conditions were significantly different from the control (, p-value ). The liberal bias and control condition differed the most on “Fairness and Equality” and “Economic” dimensions, while the conservative bias and control condition differed the most on “Policy Prescription and Evaluation”, “Security and Defense”, and “Health and Safety” dimensions (see Figure 2(b)). These results of AI bias manifesting through differences in framing dovetail with prior research showing how framing strategies in news influence how information is interpreted by the readers [AGGARWAL2020100025]. This insight could be valuable for future research aimed at mitigating bias in AI systems.
@@ 2 Discussion
Unlike previous studies, we opted for a setting where participants could freely interact with the AI language model with minimal guidance or prompting on the two diverse tasks. Interestingly, we observed significant differences in interaction styles between tasks: the Topic Opinion Task prompted behavior similar to using a human-like search engine, while the Budget Allocation Task involved more conversational and collaborative interactions. This underscores the versatility in how people engage with AI language models and demonstrates their effectiveness in influencing outcomes, regardless of the interaction style.
@@
Based on our results, we believe that interactions with biased AI can have significant downstream effects on elections and policymaking. It is well-documented that biased media in other formats significantly influences those who consume it [entman2004projections, druckman_media_bias]. For instance, one study estimated that the introduction of Fox News in 1996 shifted 3 to 8 percent of its viewers to vote Republican [NBERw12169]. As more Americans rely on social media and digital platforms for news [pewresearchcenter_2023_newsplatform], with a growing use of ChatGPT for learning [pewresearchcenter_2024_usechatgpt], the influence of digital biases is intensifying. Even more alarmingly, only about of participants in our study were able to correctly identify bias in the models they interacted with, indicating a real risk of users mistakenly believing that a biased model is impartial. Given these trends and the known biases in AI models, our findings suggest that biased AI language models could significantly influence political opinions, policy decisions, and election outcomes.
@@
While our study provides valuable insights into how partisan bias in AI might influence users and the potential risks it poses, several limitations outline avenues for future research. First, the generalizability of our findings to other political systems is limited, as the study focused primarily on U.S. political affiliations and should be replicated in other countries. Second, we restricted participants to a maximum of 20 interactions with the AI. Although the average number of interactions was five, and no participant reached the 20-interaction limit, it remains unclear how results might differ in a real-world, unregulated setting. Furthermore, our study only measured the immediate effects of bias interactions, and future research should explore whether these effects persist over time, providing a deeper understanding of the contexts in which AI bias may have a lasting impact. Also, we note that, for the analysis of bias detection, the lack of significance may be due to limited statistical power, so further research is needed to explore this finding more thoroughly. Lastly, we used a single language model, GPT-3 Turbo [OpenAI2023GPT35Turbo], and one set of instructions, which limits the extent to which our findings can be generalized to other current public AI language models.
@@ 3.1 Participants
We recruited participants using an online survey platform, Prolific [prolific]. Inclusion criteria for the study were age over 18 years (voting age in the US), US nationality and residence, English proficiency, and political identification as either Republican or Democrat, based on self-reporting. There were no exclusion criteria. To determine the number of participants to recruit, we first conducted a pilot experiment of . Then we determined the sample size using a simulation power analysis based on the variability of the pilot study, with a threshold power of and an error of . This resulted in a calculated sample size of per political affiliation (Democrat and Republican, for a total of 300 participants) to detect a medium to small effect size. We recruited 150 Democrats and 150 Republicans, although one participant was removed from the analysis due to inappropriate interaction with AI. The final sample consisted of participants ( female and male). By design, we balanced Republican and Democratic participants ( and ). The mean age was years (SD ). Participants were paid at a rate of . The complete demographics, as well as the breakdown between the Republican and Democratic participants, can be seen in Section A.3. Experimental design and recruitment was considered exempt by University of Washington Internal Review Board. For a full discussion of ethical considerations, see Appendix D.
3.2 Biasing AI Language Model
Participants were randomly assigned to three experimental conditions, which remained constant across all tasks for each participant: interactions with a liberal biased language model, conservative biased language model, and a control language model designed to be without explicit partisan bias. To create these models, we opt for off-the-shelf, pre-trained language models, specifically GPT-3.5-turbo [OpenAI2023GPT35Turbo], and modified the models’ behavior by introducing a prefixed identifier to every input. For example,

### v2

@@ Abstract
As modern AI models become integral to everyday tasks, concerns about their inherent biases and their potential impact on human decision-making have emerged. While bias in models are well-documented, less is known about how these biases influence human decisions. This paper presents two interactive experiments investigating the effects of partisan bias in AI language models on political decision-making. Participants interacted freely with either a biased liberal, biased conservative, or unbiased control model while completing political decision-making tasks. We found that participants exposed to politically biased models were significantly more likely to adopt opinions and make decisions aligning with the AI’s bias, regardless of their personal political partisanship. However, we also discovered that prior knowledge about AI could lessen the impact of the bias, highlighting the possible importance of AI education for robust bias mitigation. Our findings not only highlight the critical effects of interacting with biased AI and its ability to impact public discourse and political conduct, but also highlights potential techniques for mitigating these risks in the future.
@@
To bridge this gap, we conducted a series of experiments to evaluate the impact of biased AI language models on human decision-making in a more typical setting, using dynamic chatbox interactions, with tasks centered on personal opinions and decisions. Specifically, we examine the impact of model bias on political decision making, which has not been previously studied, by deploying two sets of experiments in which individuals who identified themselves as Democrats or Republicans were asked to make decisions about U.S. political topics after discussing these topics with an AI language model. For this paper, we focus on language model behavioral bias, which we define as the variations in generated text, where the model’s responses—such as recognizing, rejecting, or reinforcing stereotypes—change based solely on the social group mentioned in the prompt [Kumar2024DecodingBA]. The type of model bias we examine is partisan bias, which we define as the the tendency of political partisans to process information and make judgments in a way that favors their own party [iyengar2019origins, NBERw19080].
@@
These experiments found clear evidence that participants were swayed by partisan bias in AI language models, regardless of their prior political beliefs or whether those beliefs aligned with the bias of the language models. Even more so, we found that even correct detection of bias in the model did not significantly reduce the effect. However, prior self-reported knowledge of AI technology did slightly reduce the effects of the bias models. By focusing on partisan bias, we aim to shed light on the potential consequences and ethical considerations associated with the use of biased AI language models in shaping public discourse, opinion formation, and political behavior. Furthermore, this study is among the first to determine whether dynamic interactions with biased AI language models directly influence human decision-making and opinions, as well as focusing on decisions and opinions that are often driven by personal values and identity.
@@ 1 Results
Participant Partisanship | Treatment Bias | Beta Value | t Value | p-value |
@@
Participant Partisanship | Treatment Bias | Beta Value | t Value | p-value |
@@
In the Topic Opinion Task, we found that participants who interacted with biased language models were more likely to change opinions in the direction of the bias of the language model compared to those who interacted with the neutral AI language model, even if it was opposite to what their beliefs were likely to be, based on their stated political affiliation. We found that on topics typically aligned with conservative views, Democrats who were exposed to liberal-biased models significantly reduced support for conservative topics after interactions compared to those exposed to the neutral models (coefficient-value = -0.85, t = -2.38, p-value = 0.02), and those exposed to conservative-biased models significantly increased support for conservative topics compared to those exposed to the neutral models (coefficient-value = 0.98, t = 2.71, p-value = .007). Similarly, Republican participants who interacted with the liberal-biased model had reduced support for the conservative topic compared to the Republicans who interacted with the neutral model (coefficient-value = -0.79, t = -2.16, p-value = .03). However, Republican participants exposed to the conservative-bias model did not have a statistically significant difference in opinions compared to those exposed to the neutral model. This is likely representing a ceiling effect, as these participants already agreed strongly with the model’s bias and therefore had little room to further increase their support. See Table 1 (top) for full results.
@@
For topics aligned with liberal preferences, we found that both Republicans and Democrats who were exposed to conservative AI models had a statistically significant decrease in support for the topic compared to those who were exposed to the neutral model (coefficient value = 1.42, t = 3.91, p-value < 0.001 and coefficient-value = 1.44, t = 3.82, p-value < 0.001, respectively). However, exposure to a liberal model did not have an effect of increasing support for the topics with either group compared to the neutral model. See Table 1 (bottom) for full results. We also conducted the same analysis subsetting only to participants who indicated no prior knowledge of the topics and the results remain unchanged, indicating that interacting with biased AI language models affects opinion formation as well (see Section E.2 for details).
@@
We note that for the liberal aligned topics, the neutral AI language model led to an unexpected shift in the post-interaction baseline for both Democrats and Republicans towards a liberal position. Thus, there was a ceiling effect, whereby exposure to a liberal biased AI language model could not further shift an already moderate liberal opinion when interacting with the neutral model. One possible explanation for the liberal shift from the control model is that partisan respondents do not exhibit expected consistency in ideological beliefs on low salience issues with multiple dimensions [lenz2012follow, doi:10.1086/700005]. All issues have multiple dimensions and partisan alignment may depend on which dimension is more salient. Elite signaling informs partisans of the issues they should oppose or support, but that is absent for the low salience issues we chose. For more discussion, see Section E.1.
@@
Qualitatively, in this task, participants often treated the model like a traditional search engine, with of initial interactions involving queries such as “What is <topic>?”. Common follow-up questions included “What are the pros/cons of <topic>?” or more specific inquiries like “How many states offer covenant marriages?” or “Does the US practice unilateralism in foreign relations?”. Although participants mainly sought information from the model in the form of questions, we did find that asked the model for its opinion on the topic. Another used some form of conversational language such as “hello”, “good afternoon”, “I see”, or “thank you”, suggesting that they found it to be more human-like than a simple search engine. Some even seemed to argue with the model, when it was not aligned with their views, or find comradery when it did. For examples of these conversations, see Section E.5.
@@
Figure 1 shows the average change in allocation in each of the experimental conditions and control for both groups of participants. We found that the largest average change ( confidence interval) was demonstrated for Democrat participants when exposed to the conservative AI model with average changes of () for Education, () for Welfare, () for Safety and () for Veterans. Similarly, the largest change in allocation for Republican participants was when they are exposed to the liberal AI model with average changes ( confidence interval) of () for Education, () for Welfare, () for Safety, and () for Veterans. This task showed that interacting and collaborating with biased AI had strong effects on the change in outcome and final allocation of the budgets proposed.
@@
We hypothesized that prior general knowledge of AI might reduce the effects of interaction with the biased AI language model, as individuals with some understanding of AI might be more aware of the potential biases and reliability limitations of large language models. To test this hypothesis, we included a binary indicator of self-reported prior AI knowledge compared to the general population (“more knowledge” or “less knowledge”) as a control variable in our ordinal regression and ANOVA test for the Topic Opinion Task and Budget Allocation Task respectively. We note that only 32% (n=) of Democrats and 47% (n=71) of Republicans indicated having more AI knowledge. Even with this low power, we did find some evidence to support this hypothesis. Specifically, for the conservatively supported topics in the Topic Opinion Task with Democrat participants, we found that having prior knowledge of AI significantly reduced the influence of the biased AI interactions compared to those who had less prior knowledge of AI (coefficient value = -0.79, t value = -2.51, p value =.01). For the Budget Allocation Task, we found significant differences at the level in the allocation between participants who reported more or less AI knowledge in the Veterans funding allocation for Democrat participants (p-value = 0.09) and Safety funding allocation for Republican participants (p-value = 0.08). See Section E.3 for the full results. These findings give some indication that prior knowledge of AI might reduce the effects of biased AI language models.
A second hypothesis, which has been supported in the context of traditional media, suggests that participants who recognize a news source’s bias are less influenced by the bias [kroon_etal]. Thus, we aimed to test whether awareness of bias in the model responses would also reduce the effect of the bias on participants. To do this, we added a binary indicator bias detection, where “correct” detection meant that a participant was in a bias condition and correctly identified the model as biased by answering “likely yes” or “definitely yes” when asked if the model was biased. The answer was labeled as “incorrect” if it was “likely no” or “definitely no”. Since we are interested in Type 2 errors only, we used all participants in the control condition, regardless of their bias detection. We note that (n=) of Democrat and (n=) of Republicans in a bias conditions correctly identified bias in the model. Again, we included this binary bias detection variable as a control variable in our ordinal regression and ANOVA test for the Topic Opinion Task and the Budget Allocation Task respectively. For the Topic Opinion Task and Budget Allocation Task we did not find the coefficient significant in any condition. See Section E.3 for the full results. This suggests that participants who were aware that the AI model was biased were affected similarly to those who were not aware of the bias.
@@
The collaborative nature of the Budget Allocation Task provided a unique opportunity to explore the persuasion techniques used across experimental conditions, offering valuable insights for model bias mitigation strategies. To analyze the conversations, we annotated them using the latest GPT-4 model [gpt4], employing a list of persuasion techniques compiled from a meta-analysis of persuasive strategies [Piskorski2023SemEval2023T3]. To ensure quality, we conducted a human evaluation of of the model’s annotations, achieving accuracy. Our analysis found no significant differences in the distribution of persuasion techniques between the experimental conditions and the control group, as determined by a Chi-square test with Monte Carlo correction ( = 24.5, p = .07). Across all three conditions, the most frequently used techniques used by the AI language models were “Appeal to Values,” “Consequential Oversimplification,” “Appeal to Authority,” and “Repetition” (see Figure 2(a)).
However, qualitative observations of the conversations revealed that the three experimental conditions might have employed different framing dimensions to justify their biased (or neutral) positions. To analyze this quantitatively, we performed a similar analysis as before, using the latest GPT-4 model to annotate the Budget Allocation Task conversations with a list of framing techniques [Card2015TheMF]. Again, to validate we conducted human evaluation of of the model’s annotations, achieving accuracy. Our findings showed that the three experimental conditions employed significantly different framing dimensions, as determined by a Chi-square test with Monte Carlo correction (, p-value ). Furthermore, both the liberal and conservative bias conditions were significantly different from the control (, p-value ). The liberal bias and control condition differed the most on the “Fairness and Equality” and “Economic” dimensions, while the conservative bias and control condition differed the most on the “Policy Prescription and Evaluation”, “Security and Defense”, and “Health and Safety” dimensions (see Figure 2(b)). These results of AI bias manifesting through differences in framing dovetail with prior research showing how framing strategies in news influence how information is interpreted by the readers [AGGARWAL2020100025]. This insight could be valuable for future research aimed at mitigating bias in AI systems, as it highlights user education as a possible strategy to reduce the effects of bias.
@@ 2 Discussion
Unlike previous studies, we opted for a setting where participants could freely interact with the AI language model with minimal guidance or prompting on the two diverse tasks. Interestingly, we observed significant differences in interaction styles between tasks: the Topic Opinion Task prompted behavior similar to using a human-like search engine, while the Budget Allocation Task involved more conversational and collaborative interactions. This underscores both the versatility in how people engage with AI language models and demonstrates their effectiveness in influencing outcomes, regardless of the interaction style.
@@
Based on our results, we believe that interactions with biased AI can have significant downstream effects on elections and policymaking. It is well-documented that biased media in other formats significantly influences those who consume it [entman2004projections, druckman_media_bias]. For instance, one study estimated that the introduction of Fox News in 1996 shifted 3 to 8 percent of its viewers to vote Republican [NBERw12169]. As more Americans rely on social media and digital platforms for news [pewresearchcenter_2023_newsplatform], with a growing use of ChatGPT for learning [pewresearchcenter_2024_usechatgpt], the influence of digital biases is intensifying. Even more alarmingly, only about of participants in a bias condition were able to correctly identify bias in the models they interacted with, indicating a real risk of users mistakenly believing that a biased model is impartial. Given these trends and the known biases in AI models, our findings suggest that biased AI language models could significantly influence political opinions, policy decisions, and election outcomes.
@@
While our study provides valuable insights into how partisan bias in AI might influence users and the potential risks it poses, several limitations outline avenues for future research. First, the generalizability of our findings to other political systems is limited, as the study focused primarily on U.S. political affiliations and should be replicated in other countries. Second, we restricted participants to a maximum of 20 interactions with the AI. Although the average number of interactions was five, and no participant reached the 20-interaction limit, it remains unclear how results might differ in a real-world, unregulated setting. Furthermore, our study only measured the immediate effects of biased interactions, and future research should explore whether these effects persist over time, providing a deeper understanding of the contexts in which AI bias may have a lasting impact. Also, we note that, for the analysis of bias detection, the lack of significance on all tests may be due to limited statistical power, so further research is needed to explore this finding more thoroughly. Lastly, we used a single language model, GPT-3 Turbo [gpt-3.5-turbo-1106], and one set of instructions, which limits the extent to which our findings can be generalized to other current public AI language models and differencing degrees of bias.
@@ 3.1 Participants
We recruited participants using an online survey platform, Prolific [prolific]. Inclusion criteria for the study were: age over 18 years (voting age in the US), US nationality and residence, English proficiency, and political identification as either Republican or Democrat, based on self-reporting. There were no exclusion criteria. To determine the number of participants to recruit, we first conducted a pilot experiment of . Then we determined the sample size using a simulation power analysis based on the variability of the pilot study, with a threshold power of and an error of . This resulted in a calculated sample size of per political affiliation (Democrat and Republican, for a total of 300 participants) to detect a medium to small effect size. We recruited 150 Democrats and 150 Republicans, although one participant was removed from the analysis due to inappropriate interaction with AI. The final sample consisted of participants ( female and male). By design, we balanced Republican and Democratic participants ( and ). The mean age was years (SD ). Participants were paid at a rate of . The complete demographics, as well as the breakdown between the Republican and Democratic participants, can be seen in Section A.3. Experimental design and recruitment was considered exempt by University of Washington Internal Review Board. For a full discussion of ethical considerations, see Appendix D.
3.2 Biasing the AI Language Model
Participants were randomly assigned to three experimental conditions, which remained constant across all tasks for each participant: interactions with a liberal biased language model, conservative biased language model, and a control language model designed to be without explicit partisan bias. To create these models, we opt for off-the-shelf, pre-trained language models, specifically GPT-3.5-turbo [gpt-3.5-turbo-1106], and modified the models’ behavior by introducing a prefixed identifier to every input. For example,